Solve the equation by factoring. 3x2 + 13x - 10 = 0 o lğ,5} otį,53 of-53 olã, -5
Answer #1

3x? +13% -10-0 38² +15x - QX-1050 13x=15- 2x 3x² + 15x-2x-1020 3x (x+5)-2 (x+5) (3x-2)(x+5)=0 3x-220 OR *+550 X22 - 2 or x=-1pls like the answer